# See More Details: Efficient Image Super-Resolution by Experts Mining

### v1

@@ Abstract
Reconstructing high-resolution (HR) images from low-resolution (LR) inputs poses a significant challenge in image super-resolution (SR). While recent approaches have demonstrated the efficacy of intricate operations customized for various objectives, the straightforward stacking of these disparate operations can result in a substantial computational burden, hampering their practical utility. In response, we introduce SeemoRe, an efficient SR model employing expert mining. Our approach strategically incorporates experts at different levels, adopting a collaborative methodology. At the macro scale, our experts address rank-wise and spatial-wise informative features, providing a holistic understanding. Subsequently, the model delves into the subtleties of rank choice by leveraging a mixture of low-rank experts. By tapping into experts specialized in distinct key factors crucial for accurate SR, our model excels in uncovering intricate intra-feature details. This collaborative approach is reminiscent of the concept of “see more”, allowing our model to achieve an optimal performance with minimal computational costs in efficient settings. The source will be publicly made available at https://github.com/eduardzamfir/seemoredetails
@@ 2 Related Works
CNN-based SR.   In recent years, CNN-based techniques have outperformed traditional interpolation algorithms (Duchon, 1979) by learning a non-linear mapping between the input and target in an end-to-end training manner. The seminal SRCNN (Dong et al., 2014) introduced a three-layer convolutional approach for image super-resolution, later extended by works such as (Lim et al., 2017; Zhang et al., 2018b; Hui et al., 2019; Liang et al., 2021). VDSR (Kim et al., 2016) and EDSR (Lim et al., 2017) deepen networks using residual learning principles, with EDSR streamlining residual blocks for deeper training. Conversely, RCAN (Zhang et al., 2018a) introduces a novel residual-in-residual architecture for models exceeding 400 layers. While various spatial and channel attention mechanisms aim to enhance image reconstruction quality, CNN-based techniques still struggle to effectively utilize shared information across both dimensions. In this work, we aim to explore the interdependencies among the features in a computationally efficient way.
@@
Transformer-based SR.   Thanks to its remarkable performance in high-level tasks (Dosovitskiy et al., 2021), the Transformer architecture has found its way into low-level vision tasks, such as image SR. Contemporary Transformer-based approaches aim to alleviate the computational load by confining self-attention to local regions and incorporating a higher degree of locality bias into their network design. SwinIR (Liang et al., 2021) incorporates local window self-attention and a shift mechanism inspired by the Swin Transformer design(Liu et al., 2021). Meanwhile, others like ELAN (Zhang et al., 2022) or ESRT  (Lu et al., 2022) reduce the feature dimensions by splitting or down-scaling to enhance the computational efficiency. Omni-SR (Wang et al., 2023) models pixel-interactions across different axes, creating universal correlations. SRFormer (Zhou et al., 2023) optimizes the computational efficiency by employing large window self-attention through the permutation of self-attention mechanisms. However, transformer-based methods typically demand significantly higher computational resources, even with smaller model capacities.
@@
Figure 2: Architectute Overview. SeemoRe refines the feature representations via stacked Residual groups (RGs). Each RG consists of a Rank Modulating Exert (RME) and a Spatial Modulating Expert (SME). RME leverages the Mixeture of Low Rank (MoRE) to refine the global texture, while SME employs spatial enhancement experts (SEE) to supplement RME with the spatial cues.
@@
Efficiency in SR.   In recent years, the pursuit of efficient SR techniques has gained significant momentum (Li et al., 2022; Ignatov et al., 2023; Li et al., 2023; Conde et al., 2023). Consequently, researchers have introduced streamlined neural architectures (Ignatov et al., 2021), network compression (Wang et al., 2022), reparameterization (Zhang et al., 2021b), and other training strategies to cater to the demand for efficiency. Initially, efficient SR methods utilized group convolutions and cascaded block designs to boost efficiency (Ahn et al., 2018; Hui et al., 2019). Subsequent advancements introduced convolution-based spatial or channel enhancement modules (Liu et al., 2020b). More recently, ShuffleMixer (Sun et al., 2022) integrates large kernel convolutions and feature shuffling, improving both computational efficiency and high-resolution reconstruction. SAFMN (Sun et al., 2023) improves the efficiency by collecting non-local features using a shallow pyramid. Despite improvements in several efficiency aspects brought up by the aforementioned approaches, there is still scope for a better trade-off between model efficiency and the restoration performance.
@@ 3.1 Rank Modulating Expert
Mixture of Low-Rank Expertise.   As illustrated in Figure 3, from a layer normalised input tensor $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$, we use a 3$\times$3 convolution for feature projection and then we split along the channel dimension to create two distinct views $\mathbf{x}_{a}~{}\text{and}~{}\mathbf{x}_{b}\in\mathbb{R}^{H\times W\times C}$. To efficiently aggregate the pixel-wise cross-channel context, we leverage a recursive strided convolution $t$ times followed by a refinement and upsampling step, resulting in the construction of the feature pyramid denoted as $\mathbf{\hat{x}}_{b}\in\mathbb{R}^{H\times W\times C}$. The process is formulated as follows:
@@
where $\mathcal{G}(\cdot)$ and $\mathcal{E}_{i}(\cdot)$ denote the learned routing function and the output of the $i$-th expert, respectively. The sparsity inherent in the router function $\mathcal{G}(\cdot)$ optimizes computation by assigning greater weights to the top-$k$ low-rank experts. While at training time, our method learns from different experts, during inference, only the selected top-$k$ expert is utilized for computation, further enhancing the efficiency. More specifically, the inference complexity is not proportional to the number of experts. The design choices contributing towards the selection of the number of low-rank experts ($\mathcal{E}_{i}$) and the rank dimension ($R$) for memory efficient reconstruction is illustrated in Table 5 of the ablation study. We also provide the pseudocode for the proposed MoRE block in Algorithm 1.
Figure 3: Illustration of the proposed Mixture of Low-Rank Expertise (MoRE) as core block of the RME.
Algorithm 1 Mixture of Low-Rank Experise
2:  Parameters: $n$ Experts $\mathcal{E}$, Router $\mathcal{G}$, Low-Rank dimensions $R_{i}=2^{i+1}$ with $i\in\{1,...,n\}$, top-k expert $k$
3:  Compute router outputs: $\mathbf{g}=\mathcal{G}(\mathbf{\hat{x}}_{a})$
4:  Normalize weights: $\mathbf{w}=\text{Softmax}(\mathbf{g})$
5:  Select top-k expert: $w_{k}=\text{topk}(\mathbf{w},k)$
6:  if training then
7:     for each $e\in\mathcal{E}$ do
8:        $\mathbf{y}^{i}_{e}=\mathbf{W}^{3}_{R_{i}\rightarrow C}(\mathbf{W}^{1}_{C\rightarrow R_{i}}\mathbf{\hat{x}}_{a}\odot\mathbf{W}^{2}_{C\rightarrow R_{i}}\mathbf{\hat{x}}_{b})$
9:     end for
10:     Compute final output: $\mathbf{y}=\sum_{i=1}^{n}w_{i}\cdot\mathbf{y}^{i}_{e}$
11:  else
12:     Compute final output: $\mathbf{y}=w_{k}\cdot\mathbf{y}^{k}_{e}$
13:  end if
14:  Output: Final output $\mathbf{y}$
@@ 3.2 Spatial Modulating Expert
where $\odot$ is the Hadamard product, $\mathbf{W}^{4}_{C\rightarrow C}$ and $\mathbf{W}^{5}_{C\rightarrow C}$ are linear (project) layers, $\text{DConv}^{s}_{k\times k}$ denotes the striped depth-wise convolution, and $\mathbf{x}_{in}$ is the layer normalised output of the RME. The use of a large-kernel convolution facilitates a localized correlation among the pixels within the $k\times k$ window, emulating the window-based SA layers frequently employed in image restoration (Liu et al., 2021; Zamir et al., 2022; Chen et al., 2023), all the while preserving the efficiency benefits associated with convolutional layers as demonstrated in Table 4.
@@ 4 Experiments
Table 1: Comparison to efficient SR models. PSNR (dB normal-$\uparrow$) and SSIM (normal-$\uparrow$) metrics are reported on the Y-channel. Best and second best performances are highlighted. GMACS (G) are computed by upscaling to a $1280\times 720$ HR image. SeemoRe-T achieves state-of-the-art performance across all benchmarks with the lowest parameter count and computational demand. ‘-’ represents unreported results.
Method
Params GMACS Set5 Set14 BSD100 Urban100 Manga109
@@
×$\times 2$
Bicubic
- - 33.66 .9299 30.24 .8688 29.56 .8431 26.88 .8403 30.80 .9339
CARN-M (Ahn et al., 2018)
412K 91 37.53 .9583 33.26 .9141 31.92 .8960 31.23 .9193 - -
IMDN (Hui et al., 2019)
694K 159 38.00 .9605 33.63 .9177 32.19 .8996 32.17 .9283 38.88 .9774
PAN (Zhao et al., 2020)
261K 71 38.00 .9605 33.59 .9181 32.18 .8997 32.01 .9273 38.70 .9773
DRSAN (Park et al., 2021)
370K 86 37.99 .9606 33.57 .9177 32.16 .8999 32.10 .9279 - -
DDistill-SR (Wang et al., 2022)
414K 128 38.03 .9606 33.61 .9182 32.19 .9000 32.18 .9286 38.94 .9777
ShuffleMixer (Sun et al., 2022)
394K 91 38.01 .9606 33.63 .9180 32.17 .8995 31.89 .9257 38.83 .9774
SAFMN (Sun et al., 2023)
228K 52 38.00 .9605 33.54 .9177 32.16 .8995 31.84 .9256 38.71 .9771
SeemoRe-T (ours)
220K 45 38.06 .9608 33.65 .9186 32.23 .9004 32.22 .9286 39.01 .9777
×$\times 3$
Bicubic
- - 30.39 .8682 27.55 .7742 27.21 .7385 24.46 .7349 26.95 .8556
CARN-M (Ahn et al., 2018)
415K 46 33.99 .9236 30.08 .8367 28.91 .8000 27.55 .8385 - -
IMDN (Hui et al., 2019)
703K 72 34.36 .9270 30.32 .8417 29.09 .8046 28.17 .8519 33.61 .9445
PAN (Zhao et al., 2020)
261K 39 34.40 .9271 30.36 .8423 29.11 .8050 28.11 .8511 33.61 .9448
DRSAN (Park et al., 2021)
410K 43 34.41 .9272 30.27 .8413 29.08 .8056 28.19 .8529 - -
DDistill-SR (Wang et al., 2022)
414K 57 34.37 .9275 30.34 .8420 29.11 .8053 28.19 .8528 33.69 .9451
ShuffleMixer (Sun et al., 2022)
415K 42 34.40 .9272 30.37 .8423 29.12 .8051 28.08 .8498 33.69 .9448
SAFMN (Sun et al., 2023)
233K 23 34.34 .9267 30.33 .8418 29.08 .8048 27.95 .8474 33.52 .9437
SeemoRe-T (ours)
225K 20 34.46 .9276 30.44 .8445 29.15 .8063 28.27 .8538 33.92 .9460
×$\times 4$
Bicubic
- - 28.42 .8104 26.00 .7027 25.96 .6675 23.14 .6577 24.89 .7866
CARN-M (Ahn et al., 2018)
415K 33 31.92 .8903 28.42 .7762 27.44 .7304 25.62 .7694 - -
IMDN (Hui et al., 2019)
715K 41 32.21 .8948 28.58 .7811 27.56 .7353 26.04 .7838 30.46 .9075
PAN (Zhao et al., 2020)
272K 28 32.13 .8948 28.61 .7822 27.59 .7363 26.11 .7854 30.51 .9095
DRSAN (Park et al., 2021)
410K 31 32.15 .8935 28.54 .7813 27.54 .7364 26.06 .7858 - -
DDistill-SR (Wang et al., 2022)
434K 33 32.23 .8960 28.62 .7823 27.58 .7365 26.20 .7891 30.48 .9090
ShuffleMixer (Sun et al., 2022)
411K 28 32.21 .8953 28.66 .7827 27.61 .7366 26.08 .7835 30.65 .9093
SAFMN (Sun et al., 2023)
240K 14 32.18 .8948 28.60 .7813 27.58 .7359 25.97 .7809 30.43 .9063
SeemoRe-T (ours)
232K 12 32.31 .8965 28.72 .7840 27.65 .7384 26.23 .7883 30.82 .9107
Table 2: Comparison to lightweight SR Transformers. SeemoRe-L outperforms or achieves comparable performance to compared Transformers while being more efficient. Best and second best performances are highlighted. ×$\times 3$ results are in the Supplemental.
Method
Params GMACS Set5 Set14 BSD100 Urban100 Manga109
@@
×$\times 2$
Bicubic
- - 33.66 .9299 30.24 .8688 29.56 .8431 26.88 .8403 30.80 .9339
SwinIR-Light (Liang et al., 2021)
910K 244 38.14 .9611 33.86 .9206 32.31 .9012 32.76 .9340 39.12 .9783
ELAN-Light (Zhang et al., 2022)
621K 201 38.17 .9611 33.94 .9207 32.30 .9012 32.76 .9340 39.11 .9782
SRFormer-Light (Zhou et al., 2023)
853K 236 38.23 .9613 33.94 .9209 32.36 .9019 32.91 .9353 39.28 .9785
DAT-Light (Chen et al., 2023)
553K 194 38.24 .9614 34.01 .9214 32.34 .9019 32.89 .9346 39.49 .9788
SeemoRe-L (ours)
931K 197 38.27 .9616 34.01 .9210 32.35 .9018 32.87 .9344 39.49 .9790
×$\times 4$
Bicubic
- - 28.42 .8104 26.00 .7027 25.96 .6675 23.14 .6577 24.89 .7866
SwinIR-Light (Liang et al., 2021)
897K 64 32.44 .8976 28.77 .7858 27.69 .7406 26.47 .7980 30.91 .9151
ELAN-Light (Zhang et al., 2022)
601K 54 32.43 .8975 28.78 .7858 27.69 .7406 26.54 .7982 30.92 .9150
SRFormer-Light (Zhou et al., 2023)
873K 63 32.51 .8988 28.82 .7872 27.73 .7422 26.67 .8032 31.17 .9165
DAT-Light (Chen et al., 2023)
573K 50 32.57 .8991 28.87 .7879 27.74 .7428 26.64 .8033 31.37 .9178
SeemoRe-L (ours)
969K 50 32.51 .8990 28.92 .7888 27.78 .7428 26.79 .8046 31.48 .9181
Refer to caption Refer to caption
@@
HR Crop
DDistill-SR
ShuffleMixer
SwinIR-Light
DAT-Light
SeemoRe-L
@@
Table 3: Complexity Analysis. Runtime (ms, normal-$\downarrow$) and memory consumption (M, normal-$\downarrow$) averaged across 200 samples.
Method
Input Scale Time GPU Memory
DAT-Light
$[320,180]$$\times 4$ 210.12 8715.1
SwinIR-Light
131.25 6175.3
ShuffleMixer
7.40 1380.7
DDistill-SR
11.20 2822.1
SeemoRe-T (ours)
5.66 2744.9
@@
Scale
Method
Params. GMACS Urban100 Manga109
×$\times 2$

### v2

@@ Abstract
Reconstructing high-resolution (HR) images from low-resolution (LR) inputs poses a significant challenge in image super-resolution (SR). While recent approaches have demonstrated the efficacy of intricate operations customized for various objectives, the straightforward stacking of these disparate operations can result in a substantial computational burden, hampering their practical utility. In response, we introduce SeemoRe, an efficient SR model employing expert mining. Our approach strategically incorporates experts at different levels, adopting a collaborative methodology. At the macro scale, our experts address rank-wise and spatial-wise informative features, providing a holistic understanding. Subsequently, the model delves into the subtleties of rank choice by leveraging a mixture of low-rank experts. By tapping into experts specialized in distinct key factors crucial for accurate SR, our model excels in uncovering intricate intra-feature details. This collaborative approach is reminiscent of the concept of “see more”, allowing our model to achieve an optimal performance with minimal computational costs in efficient settings. The source codes will be publicly made available at https://github.com/eduardzamfir/seemoredetails
@@ 2 Related Works
In recent years, CNN-based techniques have outperformed traditional interpolation algorithms (Duchon, 1979) by learning a non-linear mapping between the input and target in an end-to-end training manner. The seminal SRCNN (Dong et al., 2014) introduced a three-layer convolutional approach for image super-resolution, later extended by works such as (Lim et al., 2017; Zhang et al., 2018b; Hui et al., 2019; Liang et al., 2021). VDSR (Kim et al., 2016) and EDSR (Lim et al., 2017) deepen networks using residual learning principles, with EDSR streamlining residual blocks for deeper training. Conversely, RCAN (Zhang et al., 2018a) introduces a novel residual-in-residual architecture for models exceeding 400 layers. While various spatial and channel attention mechanisms aim to enhance image reconstruction quality, CNN-based techniques still struggle to effectively utilize shared information across both dimensions. In this work, we aim to explore the interdependencies among the features in a computationally efficient way.
Transformer-based SR.
@@
Thanks to its remarkable performance in high-level tasks (Dosovitskiy et al., 2021), the Transformer architecture has found its way into low-level vision tasks, such as image SR. Contemporary Transformer-based approaches aim to alleviate the computational load by confining self-attention to local regions and incorporating a higher degree of locality bias into their network design. SwinIR (Liang et al., 2021) incorporates local window self-attention and a shift mechanism inspired by the Swin Transformer design(Liu et al., 2021). Meanwhile, others like ELAN (Zhang et al., 2022) or ESRT  (Lu et al., 2022) reduce the feature dimensions by splitting or down-scaling to enhance the computational efficiency. Omni-SR (Wang et al., 2023) models pixel-interactions across different axes, creating universal correlations. SRFormer (Zhou et al., 2023) optimizes the computational efficiency by employing large window self-attention through the permutation of self-attention mechanisms. However, transformer-based methods typically demand significantly higher computational resources, even with smaller model capacities.
@@
Figure 2: Architecture Overview. SeemoRe refines the feature representations via stacked Residual groups (RGs). Each RG consists of a Rank Modulating Exert (RME) and a Spatial Modulating Expert (SME). RME leverages the Mixture of Low Rank Expertise (MoRE) to refine the global texture, while SME employs spatial enhancement experts (SEE) to supplement RME with spatial cues.
Efficiency in SR.
@@
In recent years, the pursuit of efficient SR techniques has gained significant momentum (Li et al., 2022; Ignatov et al., 2023; Li et al., 2023; Conde et al., 2023). Consequently, researchers have introduced streamlined neural architectures (Ignatov et al., 2021), network compression (Wang et al., 2022), reparameterization (Zhang et al., 2021b), and other training strategies to cater to the demand for efficiency. Initially, efficient SR methods utilized group convolutions and cascaded block designs to boost efficiency (Ahn et al., 2018; Hui et al., 2019). Subsequent advancements introduced convolution-based spatial or channel enhancement modules (Liu et al., 2020b). More recently, ShuffleMixer (Sun et al., 2022) integrates large kernel convolutions and feature shuffling, improving both computational efficiency and high-resolution reconstruction. SAFMN (Sun et al., 2023) improves the efficiency by collecting non-local features using a shallow pyramid. Despite improvements in several efficiency aspects brought up by the aforementioned approaches, there is still scope for a better trade-off between model efficiency and the restoration performance.
Dynamic Networks.
Dynamic networks have been extensively studied to optimize the balance between speed and performance across various tasks. Early research employed conditional computation to selectively activate network segments at different times (Bengio et al., 2013). More recently, Mixture-of-Experts (MoE) approaches with routing architecture (Shazeer et al., 2017; Riquelme et al., 2021; Puigcerver et al., 2024) have expanded model capacity without significantly increasing inference costs, primarily enhancing the feed-forward capacity of Transformers in Natural language processing (Shazeer et al., 2017) and high-level vision tasks (Riquelme et al., 2021; Puigcerver et al., 2024). A similar idea can be found in image restoration, where Path-Restore (Yu et al., 2021) dynamically routes image patches to different network paths based on content and distortion, leveraging a difficulty-regulated reward function. In this work, our research explores the routing concept from an architecture design perspective for image super-resolution, aiming to discover the most efficient and appropriate expert to improve the feature modeling.
@@ 3.1 Rank Modulating Expert
Mixture of Low-Rank Expertise.
As illustrated in Figure 3, from a layer normalised input tensor $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$, we use a 3$\times$3 convolution for feature projection and then we split along the channel dimension to create two distinct views $\mathbf{x}_{a}~{}\text{and}~{}\mathbf{x}_{b}\in\mathbb{R}^{H\times W\times C}$. To efficiently aggregate the pixel-wise cross-channel context, we leverage a recursive strided convolution $t$ times followed by a refinement and upsampling step, resulting in the construction of the feature pyramid denoted as $\mathbf{\hat{x}}_{b}\in\mathbb{R}^{H\times W\times C}$. The process is formulated as follows:
@@
Figure 3: Illustration of the proposed Mixture of Low-Rank Expertise (MoRE) as a core block of the RME.
@@
Algorithm 1 Mixture of Low-Rank Experise
2:  Parameters: $n$ Experts $\mathcal{E}$, Router $\mathcal{G}$, Low-Rank dimensions $R_{i}=2^{i+1}$ with $i\in\{1,...,n\}$ , top-1 expert $k=1$
3:  Compute router outputs: $\mathbf{g}=\mathcal{G}(\mathbf{\hat{x}}_{a})$
4:  Normalize weights: $\mathbf{w}=\text{Softmax}(\mathbf{g})$
5:  Select top-1 expert: $w_{\text{top-1}}=\text{topk}(\mathbf{w},k=1)$
[…]
7:  if training then
8:     for each $e\in\mathcal{E}$ do
9:        $\mathbf{y}^{i}_{e}=\mathbf{W}^{3}_{R_{i}\rightarrow C}(\mathbf{W}^{1}_{C\rightarrow R_{i}}\mathbf{\hat{x}}_{a}\odot\mathbf{W}^{2}_{C\rightarrow R_{i}}\mathbf{\hat{x}}_{b})$
10:     end for
11:     Compute final output: $\mathbf{y}=\sum_{i=1}^{n}w_{i}\cdot\mathbf{y}^{i}_{e}$
12:  else
13:     Compute final output: $\mathbf{y}=w_{\text{top-1}}\cdot\mathbf{y}^{\text{top-1}}_{e}$
14:  end if
15:  Output: Final output $\mathbf{y}$
@@
where $\mathcal{G}(\cdot)$ and $\mathcal{E}_{i}(\cdot)$ denote the learned routing function and the output of the $i$-th expert, respectively. The sparsity inherent in the router function $\mathcal{G}(\cdot)$ optimizes computation by assigning greater weights to the top-$k$ low-rank experts. While at training time, our method learns from different experts, during inference, only the selected top-$k$ expert is utilized for computation, further enhancing the efficiency. More specifically, the inference complexity is not proportional to the number of experts.
Adhering to the MoE concept with $k>1$, our routing function for optimal low-rank representation extends sparse routing principles (Shazeer et al., 2017) by selecting only the top-$1$ expert. As our work is pioneering in this domain, we emphasize a more interpretable top-1 design, as shown in Figure 5(b), which allows us to streamline the model architecture and computational process, creating an efficient yet powerful image super-resolution model. Technically, both training and inference leverage dynamic expert selection based on input and model depth; however, only the top-$1$ expert per layer is utilized, with contributions from other experts weighted at zero. During inference, inactive experts are disregarded to efficiently exploit contextual information using the optimal input-dependent expert chosen by the router. This ensures consistency between training and inference, as only one expert per layer remains active, thereby mitigating potential discrepancies. In Table 9 found in the supplementary, we show that augmenting the number of top-$k$ experts can slightly improve the performance, at the cost of increased computational complexity. We hope that our network can serve as a fresh baseline for future development.
Additionally, the design choices contributing towards the selection of the number of low-rank experts ($\mathcal{E}_{i}$) and the rank dimension ($R$) for memory-efficient reconstruction is illustrated in Table 5(a) of the ablation study. We also provide the pseudocode for the proposed MoRE block in Algorithm 1. In addition to the primary analyses presented in the main text, the supplementary material offer further insights and experiments that substantiate the design decisions of our proposed MoRE module. For detailed information, refer to Tables 11 and 14.
@@ 3.2 Spatial Modulating Expert
where $\odot$ is the Hadamard product, $\mathbf{W}^{4}_{C\rightarrow C}$ and $\mathbf{W}^{5}_{C\rightarrow C}$ are linear (project) layers, $\text{DConv}^{s}_{k\times k}$ denotes the striped depth-wise convolution, and $\mathbf{x}_{in}$ is the layer normalised output of the RME. The use of a large-kernel convolution facilitates a localized correlation among the pixels within the $k\times k$ window, emulating the window-based SA layers frequently employed in image restoration (Liu et al., 2021; Zamir et al., 2022; Chen et al., 2023), all the while preserving the efficiency benefits associated with convolutional layers as demonstrated in Table 4(a).
@@ 4 Experiments
Table 1: Comparison to efficient SR models. PSNR (dB $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the Y-channel. Best and second best performances are highlighted. GMACS (G) are computed by upscaling to a $1280\times 720$ HR image. SeemoRe-T achieves state-of-the-art performance across all benchmarks with the lowest parameter count and computational demand. ‘-’ represents unreported results.
Method Params GMACS Set5 Set14 BSD100 Urban100 Manga109
@@
×$\times 2$ Bicubic - - 33.66 .9299 30.24 .8688 29.56 .8431 26.88 .8403 30.80 .9339
CARN-M (Ahn et al., 2018) 412K 91 37.53 .9583 33.26 .9141 31.92 .8960 31.23 .9193 - -
IMDN (Hui et al., 2019) 694K 159 38.00 .9605 33.63 .9177 32.19 .8996 32.17 .9283 38.88 .9774
PAN (Zhao et al., 2020) 261K 71 38.00 .9605 33.59 .9181 32.18 .8997 32.01 .9273 38.70 .9773
DRSAN (Park et al., 2021) 370K 86 37.99 .9606 33.57 .9177 32.16 .8999 32.10 .9279 - -
DDistill-SR (Wang et al., 2022) 414K 128 38.03 .9606 33.61 .9182 32.19 .9000 32.18 .9286 38.94 .9777
ShuffleMixer (Sun et al., 2022) 394K 91 38.01 .9606 33.63 .9180 32.17 .8995 31.89 .9257 38.83 .9774
SAFMN (Sun et al., 2023) 228K 52 38.00 .9605 33.54 .9177 32.16 .8995 31.84 .9256 38.71 .9771
SeemoRe-T (ours) 220K 45 38.06 .9608 33.65 .9186 32.23 .9004 32.22 .9286 39.01 .9777
×$\times 3$ Bicubic - - 30.39 .8682 27.55 .7742 27.21 .7385 24.46 .7349 26.95 .8556
CARN-M (Ahn et al., 2018) 415K 46 33.99 .9236 30.08 .8367 28.91 .8000 27.55 .8385 - -
IMDN (Hui et al., 2019) 703K 72 34.36 .9270 30.32 .8417 29.09 .8046 28.17 .8519 33.61 .9445
PAN (Zhao et al., 2020) 261K 39 34.40 .9271 30.36 .8423 29.11 .8050 28.11 .8511 33.61 .9448
DRSAN (Park et al., 2021) 410K 43 34.41 .9272 30.27 .8413 29.08 .8056 28.19 .8529 - -
DDistill-SR (Wang et al., 2022) 414K 57 34.37 .9275 30.34 .8420 29.11 .8053 28.19 .8528 33.69 .9451
ShuffleMixer (Sun et al., 2022) 415K 42 34.40 .9272 30.37 .8423 29.12 .8051 28.08 .8498 33.69 .9448
SAFMN (Sun et al., 2023) 233K 23 34.34 .9267 30.33 .8418 29.08 .8048 27.95 .8474 33.52 .9437
SeemoRe-T (ours) 225K 20 34.46 .9276 30.44 .8445 29.15 .8063 28.27 .8538 33.92 .9460
×$\times 4$ Bicubic - - 28.42 .8104 26.00 .7027 25.96 .6675 23.14 .6577 24.89 .7866
CARN-M (Ahn et al., 2018) 415K 33 31.92 .8903 28.42 .7762 27.44 .7304 25.62 .7694 - -
IMDN (Hui et al., 2019) 715K 41 32.21 .8948 28.58 .7811 27.56 .7353 26.04 .7838 30.46 .9075
PAN (Zhao et al., 2020) 272K 28 32.13 .8948 28.61 .7822 27.59 .7363 26.11 .7854 30.51 .9095
DRSAN (Park et al., 2021) 410K 31 32.15 .8935 28.54 .7813 27.54 .7364 26.06 .7858 - -
DDistill-SR (Wang et al., 2022) 434K 33 32.23 .8960 28.62 .7823 27.58 .7365 26.20 .7891 30.48 .9090
ShuffleMixer (Sun et al., 2022) 411K 28 32.21 .8953 28.66 .7827 27.61 .7366 26.08 .7835 30.65 .9093
SAFMN (Sun et al., 2023) 240K 14 32.18 .8948 28.60 .7813 27.58 .7359 25.97 .7809 30.43 .9063
SeemoRe-T (ours) 232K 12 32.31 .8965 28.72 .7840 27.65 .7384 26.23 .7883 30.82 .9107
Table 2: Comparison to lightweight SR Transformers. PSNR (dB $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the Y-channel. Best and second best performances are highlighted. GMACS (G) are computed by upscaling to a $1280\times 720$ HR image. SeemoRe-L outperforms or achieves comparable performance to compared Transformers while being more efficient. ×$\times 3$ results are in the Supplemental.
Method Params GMACS Set5 Set14 BSD100 Urban100 Manga109
@@
×$\times 2$ Bicubic - - 33.66 .9299 30.24 .8688 29.56 .8431 26.88 .8403 30.80 .9339
SwinIR-Light (Liang et al., 2021) 910K 244 38.14 .9611 33.86 .9206 32.31 .9012 32.76 .9340 39.12 .9783
ELAN-Light (Zhang et al., 2022) 621K 201 38.17 .9611 33.94 .9207 32.30 .9012 32.76 .9340 39.11 .9782
SRFormer-Light (Zhou et al., 2023) 853K 236 38.23 .9613 33.94 .9209 32.36 .9019 32.91 .9353 39.28 .9785
ESRT (Lu et al., 2022) 751K 191 38.03 .9600 33.75 .9184 32.25 .9001 32.58 .9318 39.12 .9774
SwinIR-NG (Choi et al., 2023) 1181K 274 38.17 .9612 33.94 .9205 32.31 .9013 32.78 .9340 39.20 .9781
DAT-Light (Chen et al., 2023) 553K 194 38.24 .9614 34.01 .9214 32.34 .9019 32.89 .9346 39.49 .9788
SeemoRe-L (ours) 931K 197 38.27 .9616 34.01 .9210 32.35 .9018 32.87 .9344 39.49 .9790
×$\times 4$ Bicubic - - 28.42 .8104 26.00 .7027 25.96 .6675 23.14 .6577 24.89 .7866
SwinIR-Light (Liang et al., 2021) 897K 64 32.44 .8976 28.77 .7858 27.69 .7406 26.47 .7980 30.91 .9151
ELAN-Light (Zhang et al., 2022) 601K 54 32.43 .8975 28.78 .7858 27.69 .7406 26.54 .7982 30.92 .9150
SRFormer-Light (Zhou et al., 2023) 873K 63 32.51 .8988 28.82 .7872 27.73 .7422 26.67 .8032 31.17 .9165
ESRT (Lu et al., 2022) 751K 68 32.19 .8947 28.69 .7833 27.69 .7379 26.39 .7962 30.75 .9100
SwinIR-NG (Choi et al., 2023) 1201K 63 32.44 .8980 28.83 .7870 27.73 .7418 26.61 .8010 31.09 .9161
DAT-Light (Chen et al., 2023) 573K 50 32.57 .8991 28.87 .7879 27.74 .7428 26.64 .8033 31.37 .9178
SeemoRe-L (ours) 969K 50 32.51 .8990 28.92 .7888 27.78 .7428 26.79 .8046 31.48 .9181
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
HR Crop DDistill-SR ShuffleMixer SwinIR-Light DAT-Light SeemoRe-L
@@
Table 3: Complexity Analysis. Runtime (ms, $\downarrow$) and memory consumption (M, $\downarrow$) averaged across 200 samples using a NVIDIA RTX 4090 device.
Method Input Scale Time GPU Memory
DAT-Light $[320,180]$$\times 4$ 210.12 8715.1
SwinIR-Light 131.25 6175.3
SRFormer-Light 103.95 7270.1
SeemoRe-L (ours) 17.99 9531.6
ShuffleMixer 7.40 1380.7
DDistill-SR 11.20 2822.1
SeemoRe-T (ours) 5.66 2744.9
(a) Runtime and memory consumption.
Scale Method Params. GMACS Urban100 Manga109